OpenStudy (anonymous):

Find x and y if (x-yi)(2+3i) = x+2i/1-i.

OpenStudy (anonymous):

simple the Left hand side and rationalize the Right hand side of equation ; then compare real and imaginery parts of complex numbers from both side

OpenStudy (anonymous):

well if you dont mind can yu do after this step .. (2x + 3y) + ( 3x - 2y )i = x + 2i / 1 - i. ?

OpenStudy (anonymous):

well for rationalizing the right hand side u have \[\frac{x+2i}{1-i}=\frac{x+2i}{1-i}\frac{1+i}{1+i}=\frac{(x+2i)(1+i)}{2}\] since (1+i)(1-i)=1-i^2=1-(-1)=2

OpenStudy (anonymous):

@mukushla thank you :D

OpenStudy (anonymous):

welcome